OpenStudy (anonymous):

The cost of three notebooks and four pencils is $8.50. The cost of five notebooks and eight pencils is 14.50. Determine the cost of one notebook and the cost of one pencil.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Step 1 Read the problem carefully.Decide what unknowns are asked for?

OpenStudy (anonymous):

let notebooks be x and pencils be y

OpenStudy (anonymous):

so 3x+4y = £8.50 5x+8y= £14.50

OpenStudy (anonymous):

1. 3x+4y =8.50 2. 5x+8y =14.50 times eq1 by 2 1. 6x+8y=17 now minus eq2 form eq1 6x - 5x = x 8y- 8y= 0 17- 14.50 = 2.50 sox=2.50

OpenStudy (anonymous):

now subsitute x into eq2 5x+8y=14.50 5*2.50 = 12.50 12.50 +8y = 14.50 take away 12.50 from both sides 8y = 2 divide both sides by 8 y = 2/8 or 0.25 or 25 pence

OpenStudy (anonymous):

@skullpatrol they are asking for how much 1 notebook and pencil would cost